OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

-3 - 6 / 10 - x = -9 / 10 - x = -3 / 2 -18 = -30 + 3x 12 = 3x x=4